# Tradeoffs Between Alignment and Helpfulness in Language Models with Representation Engineering

### v3

@@ 3.1 Interpretation of results – a tradeoff between alignment and usefulness
The combination of the two results show that alignment improves linearly with the norm of the vectors injected in representation engineering while helpfulness is decreased quadratically. This means that when injecting vectors of small norms, the improvement of alignment is initially faster than the decrease in helpfulness, possibly indicating that there is a regime where prompt engineering is more effective. See figure 1 for an illustration of this intuition.
@@ 4 Empirical results
Here we will calculate alignment and helpfulness as defined above and observe how they change as we increase the vector norms of representation engineering. In principle, theorem 1 shows the dynamics of behavior flipping its signs due to representation engineering, thus to demonstrate it, we use representation engineering to show an increase in alignment of an unaligned pretrained model (specifically we use Llama 2 13B Touvron et al. (2023)), and a decrease in alignment of an aligned RLHF model (we use Llama 2 13B chat Touvron et al. (2023)). For the unaligned model, we calculate the behavior expectation w.r.t. behaviors "harmless" and "not-racist", as a function of representation engineering vector coefficients and show an improvement in alignment. For the aligned model, we do the same for the behaviors "harmful" and "racist" and show a decrease in alignment. Our experiments show an effect of representation engineering on alignment that matches theorem 1. Then, we calculate helpfulness as the probability of answering queries correctly when the model is injected with the same behavior altering vectors. Our experiments demonstrate that helpfulness changes as in theorem 2. Additional experimental details can be found in appendix E.
We follow the work of Zou et al. (2023a) to extract the vectors used in representation engineering: We use datasets comprised of pairs of positive and negative statements w.r.t. a behavior. The difference between the representations of the pairs are used to find latent space directions that can steer the model’s responses from negative to positive behaviors or vice versa. For the "harmful" behavior on the aligned model, we extracted harmful and unharmful instructions from AdvBench Robey et al. (2021; 2022) and shareGPT respectively. For "harmless" behavior on the the unaligned model, the above approach of contrasting positive and negative requests does not work, since the model did not undergo alignment, so it equally agrees to answer both types of requests. Thus, the produced engineered representations never steer the model towards not answering any request. So, inspired by the method of preference learning, we contrast aligned and misaligned responses to harmful instructions from AdvBench. For "racism" on the aligned model, we used biased and unbiased statements from the StereoSet dataset Nadeem et al. (2020). For "not-racist" on the unaligned model, we used the racist statements from above, followed by aligned and misaligned responses, and contrasted them. The vectors obtained here were used to calculate behavior expectation and helpfulness of the model as the norm of the vectors increased.
@@
To calculate harmful behavior expectation, we used the behavior scoring function that assigns an answer $B(answer)=\pm 1$ if the model agrees/refuses to answer a harmful instruction and calculated its expectation value, which amounts to the difference between probabilities of agreeing and disagreeing to answer. To calculate the racism behavior expectation, we used a behavior scoring function that assigns an answer $B(answer)=\pm 1$ to agreeing/disagreeing with a racist statement, and calculated the expectation value of this function w.r.t.  the model distribution, which amounts to the difference between probabilities of agreeing and disagreeing with a racist statement.
@@
To calculate helpfulness, we queried the model with multiple choice questions from the MMLU dataset Hendrycks et al. (2020) over a variety of domains (international law, medical genetics, high school computer science) and calculated the probability that the model assigns the correct one. This was measured as a function of injected vector coefficients inserted to the model for the behaviors above. Figure 5a(c) shows this for the harmless (harmful) behavior vectors on the misaligned (aligned) model and 5b(c) shows this for the not-racist (racist) behavior vectors for the misaligned (aligned) model. Here we restricted the probabilities to the answers of the multiple choice question, A,B,C,D.
@@
According to theorem 2 and corollary 2, around $r_{e}=0$, helpfulness should decrease parabolically, which can be seen by Taylor expanding the bounds, yielding $\mathbb{P}_{correct}\approx f(0)+\frac{1}{2}f^{\prime\prime}(0)r_{e}^{2}$ where $f^{\prime\prime}(0)<0$. Then, according to corollary 3, since there are $N=4$ answers to choose from, the probability of the correct answer should on average converge to $1/4$. To demonstrate this behavior in the empirical measurements, we plot a bound of the form of theorem 2 with the boundary conditions of corollary 3 (for further explanation on the theoretical justification of using this bound see appendix E.4). We do so with free parameter $\lambda\sigma\beta$, and $\alpha$ set to $1$ (in our theoretical bound it is smaller, but it is due to the bound being centered at $r_{e}=0$ while the peak is not guaranteed to be, hence it may overshoot), from which we find $\lambda\sigma\beta$ in the range of $0.33-0.66$. With our empirically estimated values of $\lambda\sigma\beta$ from direct measurement of the noise injected due to representation engineering, in the range $0.1-0.4$ (see details on the empirical estimation in appendix A.3).
@@
Figure 5: Helpfulness measurement: the probability assigned to the correct answer to questions from different MMLU tests (international law, medical genetics, high school computer science), as a function of representation engineering vector coefficients injected to the model. Here the probability of the correct answer was measured relative to the answers A, B, C, D. The red line plots the bound of theorem 2 for free parameters on "international law". (a) Harmless behavior expectation of Llama 2 13B as a function of coefficient of injected harmful PCA vectors. (b) Racism behavior expectation of Llama 2 13B as a function of coefficient of injected bias PCA vectors. (c) Harmful behavior expectation of Llama 2 13B as a function of coefficient of injected harmful PCA vectors. (d) Racism behavior expectation of Llama 2 13B chat as a function of coefficient of injected bias PCA vectors.
@@
In appendix E, we perform the same experiment but without restricting the answers to A,B,C,D (over the full vocabulary). The same result of theorem 2 is observed, though the helpfulness decays to zero instead of $1/4$ due to $1/|V|$ scaling of corollary 3. The average helpfulness is usually parabolic around $r_{e}=0$ but sometimes has a plateau instead, which corresponds to the helpfulness for different queries not being peaked exactly at $r_{e}=0$, for them the bound is not tight, hence a decrease that is not parabolic around $r_{e}=0$ can be observed.
@@ 5 Discussion
In principle, our framework may be generalized for theoretically analyzing the effects of normal finetuneing on alignment and helpfulness, as it too could be summed up to a change in the representations of the LLM to maximize the likelihood of desired outputs. In particular, each step in preference learning is equivalent to a representation engineering injection with coefficient that equals to the learning rate (see appendix G). However, we leave this for future work as the changes of a finetuning process to a model’s representations are different than representation engineering’s change: While finetuning creates small changes to the model’s representation at each training step on several different behaviors, that sums up to a large overall change, representation engineering takes a large step in one behavioral direction, disregarding other behavior verticals. As a result, the change to the representations in a representation engineering process on one behavior creates random noise on the others (assumption 3), unlike a finetuning process where this does not necessarily happen. Hence in regards of maintaining helpfulness, finetuning has an advantage, however, representation engineering does enjoy the benefit of an online controllable step size in the desired behavior which allows to effectively manipulate the specific behavior at inference time.
@@ A.1 Introduction of assumptions
We assume that for small coefficients of representation engineering $r_{e}$, the norm of the change to the last hidden layer representation is linear in $r_{e}$:
@@
Let $\mathbb{P}_{\theta,r_{e}}(\cdot|q)$ be a language model prompted with query $q$. The change to the last hidden layer representation due to representation engineering with coefficient $r_{e}$, denoted by $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$ satisfies:
@@
A representation of an answer to a query is defined as the latent space embedding of the answer’s token, $U^{T}e_{token}$, where $e_{i}$ is the one-hot vector of the token $i$ and $U$ is the matrix from the last layer’s hidden dimension to the vocabulary. We assume that the change to the last hidden layer of the model due to representation engineering SVM classifies representations of positive and negative answers to a query with margin $\Delta$:
Assumption 2.
Given a query $q$, the change to the last hidden layer of a model due to representation engineering, $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$ , SVM classifies the representations of positive and negative answers to a query $q$ with margin $\Delta$, where the positive and negative answers are defined with respect to a behavior scoring function $B:\Sigma\rightarrow\{-1,+1\}$:
For queries whose topic is unrelated to the behavior with respect to which representation engineering is performed, we expect the change to the last layer representation to be somewhat random, since the injected vectors take behavior related queries and shift them to positive behavior, but queries that are unrelated to behaviors can be shifted in random directions: Intuitively, the correction to the final layer representation has no preference for a specific token over the correct answer token, so an incorrect answer is just as likely to be on one side or the other of the plane defined by the vertical $\delta r_{e}(q)$ that passes through the correct answer representation.
Assumption 3.
When sampling a query $q$ that is unrelated to the behavior of representation engineering, the vector $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$, i.e., the resulting change to the last hidden layer representation due to the steering vectors from all layers, is random with the following coordinate-wise distribution:
$\langle\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|},U^{T}e_{i}\rangle\sim N(0,\sigma^{2})$ (10)
It is used in theorem 2 to formalize that representation engineering is a "perpendicular" direction to the query’s relevant answer representations.
@@ A.2 Discussion of assumptions
SVM classification with margin $\Delta$ (assumption 2): We expect the representation engineered vectors $r_{e}$ to be good classifiers because they are obtained by methods of finding directions in the latent space that maximize the distance between representations of positive and negative textual statements. For example, in Zou et al. (2023a) the first principle component is used as a steering vector, obtained via $pca_{1}=argmax_{v}\{\mathbb{E}_{good,bad}[|\langle v,r_{good}-r_{bad}\rangle|^{2}]\}$ and in Jorgensen et al. (2023) the steering vector is obtained as the average of difference between positive and negative statements $\frac{1}{N}\sum_{i=1}^{N}(r^{i}_{good}-r^{i}_{bad})$. In these examples, $r_{good}$ and $r_{bad}$ are representations of queries and not the latent space embedding of the answers, as in the definition of $\Delta$-representation-separability, but we expect the steering vectors to behave similarly on them. In subsection A.3, we show that indeed $\delta r_{e}(q)$ clusters positive and negative responses to harmful queries in the model’s latent space. In appendix F we also formulate a theorem equivalent to theorem 1, but with an imperfect classifier.
Normal token logit change (assumption 3): Random vectors in high dimension tend to have a normal distribution around a sphere. If we take an isotropic normal distribution on a sphere and look at the coordinate-wise marginal probability, then it too will look Gaussian. The change to logits does not need to be exactly normal for theorem 2 to hold, it suffices that on average the correct answer logit does not increase relative to the incorrect answers’ logits (i.e.  some incorrect answer logits are enhanced w.r.t.the correct answer logit due to representation engineering and some are reduced), but we choose this assumption due to the theoretical motivation and as can be seen in A.3, this assumption is approximately satisfied in real LLMs.
@@ A.3 Experiments for assumptions
Figure 6: Linear increase in the norm of $U\delta r_{e}(q)$ for small coefficients, when injected with "racist" vectors.
In figures 7 and 8 we plot the change in norm for Llama 2 13B chat (injected with racist vectors) and Llama 2 13B (injected with not racist vectors) respectively, on the datasets "international law", "medical genetics" and "high school computer science". We add fitted curves to estimate $\lambda$. We find that it is in the range $40-60$.
@@
Normal assumption
As proposed in assumption 3, we show here that the projection of a given answer on the representation change $\delta r_{e}(q)$ is approximately normal, with mean 0. (Assuming the question asked is not connected to the property we are changing with the representation engineering). In assumption 3 we looked at the normalized change: $\langle\frac{\delta r_{e}(q)}{\lvert\lvert\delta r_{e}(q)\rvert\rvert},U^{T}e_{i}\rangle$. Here we will look at $\langle\delta r_{e}(q),U^{T}e_{i}\rangle$, so we expect the distribution to be:
$\langle\delta r_{e}(q),U^{T}e_{i}\rangle\sim N(0,\sigma^{2}\lvert\lvert\delta r_{e}(q)\rvert\rvert^{2})$
Meaning the standard deviation scales linearly with the norm of $\delta r_{e}(q)$. Since $r_{e}$ scales linearly with $\delta r_{e}(q)$, we expect the standard deviation to also scale linearly with $r_{e}$. In practice we look at $\langle\delta r_{e}(q),U^{T}(e_{i}-e_{correct})\rangle$, the advantage of this being that it shows explicitly that the correct answer logit change is not always on one side of the Gaussian tail.
To create the plot, for each question in a dataset, we look at the top 10 answers $e_{i},i\in[10]$ (with no representation engineering). Now for a given $r_{e}$ coefficient, we calculate the projection of these answers on $\delta r_{e}(q)$. We then aggregate these projections for all the questions in a few dataset and look at their histogram and at their standard deviation. We repeat this for different $r_{e}$ norms.
@@
The tangent of the curve of figure 9c,d is $\lambda\sigma$, as the curve is the standard deviation of $\langle\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|},U^{T}e_{i}\rangle\cdot|\delta r_{e}(q)|=\langle\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|},U^{T}e_{i}\rangle\cdot\lambda r_{e}$, from assumption 1, and the inner product is a normal variable of standard deviation $\sigma$, hence the tangent is $\lambda\sigma$. From the linear curve, we estimate $\lambda\sigma=0.5$, thus $\lambda\sigma\beta\approx 0.8\cdot 0.5$, as it is the mean of a half-normal distribution with parameter $\lambda\sigma$, which is approximately $0.8\lambda\sigma$.
@@
In figure 11 and 12, we plot the distance between the centers of representation clusters for positive and negative answers to harmful queries as the norm of harmful vectors is increased, for Llama 2 13B chat and Llama 2 13B respectively. As can be seen, the distance between the clusters increases, which corresponds to an increase in $\mathbb{E}[\langle\delta r_{e}(q),U^{T}(e_{good}-e_{bad})\rangle]$. We can define a range of coefficients in which the increase is bounded from below by a linear curve of the form in equation 12, meaning that the change in the model’s representation separates the positive and negative answer representations, similarly to the definition of $\Delta$-representation separability, but with mean instead of min. Thus by equation 12, the tangent of the lower bounding lines of figures 11 and 12 are an estimate for $\Delta\lambda$. From, this we get that $\Delta\lambda$ is approximately $0.1-0.3$

### v4

@@ 3.1 Interpretation of results – a tradeoff between alignment and usefulness
The combination of the two results show that alignment improves linearly with the norm of the vectors injected in representation engineering while helpfulness is decreased quadratically. This means that when injecting vectors of small norms, the improvement of alignment is initially faster than the decrease in helpfulness, possibly indicating that there is a regime where representation engineering is more effective. See figure 1 for an illustration of this intuition.
@@ 4 Empirical results
Here we will calculate alignment and helpfulness as defined above and observe how they change as we increase the vector norms of representation engineering. In principle, theorem 1 shows the dynamics of behavior flipping its signs due to representation engineering, thus to demonstrate it, we use representation engineering to show an increase in alignment of an unaligned pretrained model (specifically we use Llama 2 13B Touvron et al. (2023)), and a decrease in alignment of an aligned RLHF model (we use Llama 2 13B chat Touvron et al. (2023)). In appendix E, we perform the experiments on Llama 3.1 8B Dubey et al. (2024) as well. For the unaligned model, we calculate the behavior expectation w.r.t. behaviors “harmless" and “not-racist", as a function of representation engineering vector coefficients and show an improvement in alignment. For the aligned model, we do the same for the behaviors “harmful" and “racist" and show a decrease in alignment. Our experiments show an effect of representation engineering on alignment that matches theorem 1. Then, we calculate helpfulness as the probability of answering queries correctly when the model is injected with the same behavior altering vectors. Our experiments demonstrate that helpfulness changes as in theorem 2. Additional experimental details can be found in appendix E.
We follow the work of Zou et al. (2023a) to extract the vectors used in representation engineering: We use datasets comprised of pairs of positive and negative statements w.r.t. a behavior. The difference between the representations of the pairs are used to find latent space directions that can steer the model’s responses from negative to positive behaviors or vice versa. For the “harmful" behavior on the aligned model, we extracted harmful and unharmful instructions from AdvBench Robey et al. (2021; 2022) and shareGPT respectively. For “harmless" behavior on the the unaligned model, the above approach of contrasting positive and negative requests does not work, since the model did not undergo alignment, so it equally agrees to answer both types of requests. Thus, the produced engineered representations never steer the model towards not answering any request. So, inspired by the method of preference learning, we contrast aligned and misaligned responses to harmful instructions from AdvBench. For “racism" on the aligned model, we used biased statements from the StereoSet dataset Nadeem et al. (2020) followed by aligned and misaligned responses, and contrasted them. For “not-racist" on the unaligned model, we flipped the sign of the vectors to reverse the steering direction. The vectors obtained here were used to calculate behavior expectation and helpfulness of the model as the norm of the vectors increased.
@@
To calculate harmful behavior expectation, we sampled full responses to harmful instructions and used the behavior scoring function that assigns an answer $B(answer)=\pm 1$ if the model answers a harmful instruction or refuses to and calculated its expectation value, which is the difference between probabilities of fulfilling and not fulfilling the instruction. To calculate the racism behavior expectation, sampled full responses to racist statements and used a behavior scoring function that assigns an answer $B(answer)=\pm 1$ to agreeing/disagreeing with a racist statement, and calculated the expectation value of this function w.r.t.  the model distribution, which is the difference in probabilities of agreeing and disagreeing with a racist statement.
@@
To calculate helpfulness, we tested the model on two tasks. The first is knowledge based question answering (MMLU), which allows a clean test for the single token theoretical results (theorem 2), and the second is coding tasks (HumanEval), which allows to verify the single token results persist for tasks with multiple-token answers. Importantly, we injected the model with the same vectors used to alter the model’s behavior in the alignment measurement. In appendix A.3 we show that while these vectors typically separate between aligned and misaligned responses in the model’s latent space when the prompt is related to the behavior in question, when prompted with knowledge based questions, the tokens enhanced are random with a trend that fits our theoretical assumptions.
@@
First, we queried the model with multiple choice questions from the MMLU dataset Hendrycks et al. (2020) over a variety of domains (international law, medical genetics, high school computer science) and calculated the probability that the model assigns the correct one answer. This was measured as a function of injected vector coefficients inserted to the model for the behaviors above. Figure 5a(c) shows this for the harmless (harmful) behavior vectors on the misaligned (aligned) model and 5b(c) shows this for the not-racist (racist) behavior vectors for the misaligned (aligned) model. Here we restricted the probabilities to the answers of the multiple choice question, A,B,C,D. We also performed the experiment in a sampling setting, where we sampled full responses to the questions, and calculated the accuracy on the dataset, where similar trends were observed (see appendix E).
@@
According to theorem 2, around $r_{e}=0$, helpfulness should decrease parabolically, which can be seen by Taylor expanding the bounds, yielding $P_{correct}\approx f(0)+\frac{1}{2}f^{\prime\prime}(0)r_{e}^{2}$ where $f^{\prime\prime}(0)<0$. Then, according to corollary 2, since there are $N=4$ answers to choose from, the probability of the correct answer should on average converge to $1/4$. To demonstrate this behavior in the empirical measurements, we plot a bound of the form of theorem 2 with the boundary conditions of corollary 2 (for further explanation on the theoretical justification of using this bound see appendix E.4). We do so with free parameter $\lambda\sigma\beta$, and $\alpha$ set to $1$ (in our theoretical bound it is smaller, but it is due to the bound being centered at $r_{e}=0$ while the peak is not guaranteed to be, hence it may overshoot), from which we find $\lambda\sigma\beta$ in the range of $0.33-0.66$. With our empirically estimated values of $\lambda\sigma\beta$ from direct measurement of the noise injected due to representation engineering, in the range $0.1-0.4$ (see details on the empirical estimation in appendix A.3).
@@
Figure 5: Helpfulness measurement: the probability assigned to the correct answer to questions from different MMLU tests (international law, medical genetics, high school computer science), as a function of representation engineering vector coefficients injected to the model. Here the probability of the correct answer was measured relative to the answers A, B, C, D. The red line plots the bound of theorem 2 for free parameters on “international law". (a) Helpfulness of Llama 2 13B with harmful PCA vectors. (b) Helpfulness of Llama 2 13B with bias PCA vectors. (c) Helpfulness of Llama 2 13B chat with harmful PCA vectors. (d) Helpfulness of Llama 2 13B chat with bias PCA vectors.
@@
Next, we tested the model’s coding skills with the humaneval dataset. as can be seen in figure 6, the model’s performance is peaked around $r_{e}=0$, and it decays parabolically ar $r_{e}$ increases. We note that the asymmetry between positive and negative coefficients is captured in our theoretical bounds.
Figure 6: Helpfulness measurement on humaneval of Llama 2 13B chat as a function of coefficient of injected harmfulness (a) and racism (b) PCA vectors.
@@ 5 Discussion
While representation engineering is an emerging field, editing interpretable features of models on the representation level in order to control them scales to SOTA models such as Sonnet’s Claude Templeton (2024). In principle, our framework may be generalized for theoretically analyzing the effects of normal finetuneing on alignment and helpfulness, as it too amounts to a change in the LLM representations to maximize the likelihood of desired outputs. In particular, each step in preference learning is equivalent to a representation injection with coefficient that equals to the learning rate (see appendix G). However, we leave this for future work, as finetuning creates small changes to the model’s representation at each training step on several behaviors, that sums to a large overall change, while representation engineering takes a large step in one behavioral direction. As a result, the change to the representations in a representation engineering process on one behavior creates random noise on the others (assumption 2), unlike a finetuning process where this does not necessarily happen. Hence in regards of maintaining helpfulness, finetuning has an advantage, however, representation engineering does enjoy the benefit of an online controllable step size in the desired behavior which allows to effectively manipulate the specific behavior at inference time.
@@ A.1 Introduction of assumptions
A representation of an answer to a query is defined as the latent space embedding of the answer’s token, $U^{T}e_{token}$, where $e_{i}$ is the one-hot vector of the token $i$ and $U$ is the matrix from the last layer’s hidden dimension to the vocabulary. We assume that the representations of positive and negative answers to a query are linearly separable, and that the change to the last hidden layer of the model due to representation engineering linearly classifies them with margin $\Delta$:
@@
Given a query $q$, the change to the last hidden layer of a model due to representation engineering, $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$ , linearly classifies the representations of positive and negative answers to a query $q$ with margin $\Delta$, where the positive and negative answers are defined with respect to a behavior scoring function $B:\Sigma\rightarrow\{-1,+1\}$:
For queries whose topic is unrelated to the behavior with respect to which representation engineering is performed, we expect the change to the last layer representation to be somewhat random on the highest probability tokens as they answer a question that is unrelated to the behavior whose vectors are injected to the model. Intuitively, the change to the final layer representation has no preference for a correct token over an incorrect token, so an incorrect answer is just as likely to be on one side or the other of the plane defined by the vertical $\delta r_{e}(q)$ that passes through the correct answer representation.
Assumption 2.
When sampling an answer to a query $q$ that is unrelated to the behavior of representation engineering, the vector $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$, i.e., the resulting change to the last hidden layer representation due to the steering vectors from all layers, is random with the following coordinate-wise distribution on the $T$ highest probability tokens making $1-\epsilon$ of the probability mass:
$\langle\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|},U^{T}e_{i}\rangle\sim D$ (8)
Where $D$ is some continuous distribution with variance $\sigma^{2}>0$.
This defines a random directionality of $\delta r_{e}(q)$ w.r.t. the representations of answers. It is used in theorem 2 to formalize that representation engineering is a “perpendicular" direction to the query’s relevant answer representations.
Finally, we assume that for small coefficients of representation engineering $r_{e}$, the norm of the change to the last hidden layer representation is linear in $r_{e}$:
Assumption 3.
Let $P_{\theta,r_{e}}(\cdot|q)$ be a language model prompted with query $q$. The change to the last hidden layer representation due to representation engineering with coefficient $r_{e}$, denoted by $\delta r_{e}(q)=r^{(L)}(q,r_{e})-r^{(L)}(q,0)$ satisfies:
@@ A.2 Discussion of assumptions
Linear classification with margin $\Delta$ (assumption 1): We expect the representation engineered vectors $r_{e}$ to be good classifiers because they are obtained by methods of finding directions in the latent space that maximize the distance between representations of positive and negative textual statements. For example, in Zou et al. (2023a) the first principle component is used as a steering vector, obtained via $pca_{1}=argmax_{v}\{\mathbb{E}_{good,bad}[|\langle v,r_{good}-r_{bad}\rangle|^{2}]\}$ and in Jorgensen et al. (2023) the steering vector is obtained as the average of difference between positive and negative statements $\frac{1}{N}\sum_{i=1}^{N}(r^{i}_{good}-r^{i}_{bad})$. In these examples, $r_{good}$ and $r_{bad}$ are representations of queries and not the latent space embedding of the answers, as in the definition of $\Delta$-representation-separability, but we expect the steering vectors to behave similarly on them. In subsection A.3, we show that indeed $\delta r_{e}(q)$ clusters positive and negative responses to harmful queries in the model’s latent space. In appendix F we also formulate a theorem equivalent to theorem 1, but with an imperfect classifier.
Random directionality of last hidden layer change (assumption 2): When answering queries that are unrelated to the behavior being enhanced by representation engineering, the directionality of the injected vectors are expected to be random w.r.t. the representations of the answers to the query. Therefore, the highest probability tokens are expected to be injected with random noise. We validate this in the next subsection, by looking at the noise injected into the top 10 highest probability tokens in knowledge queries (which typically make over 90% of the probability mass).
@@ A.3 Experiments for assumptions
Figure 7: Linear increase in the norm of $U\delta r_{e}(q)$ for small coefficients, when injected with “racist" vectors.
In figures 8 and 9 we plot the change in norm for Llama 2 13B chat (injected with racist vectors) and Llama 2 13B (injected with not racist vectors) respectively, on the datasets “international law", “medical genetics" and “high school computer science". We add fitted curves to estimate $\lambda$. We find that it is in the range $40-60$.
@@
Random logit noise assumption
As proposed in assumption 2, we show here that the projection of a given answer on the representation change $\delta r_{e}(q)$ is random. (Assuming the question asked is not connected to the property we are changing with the representation engineering). In assumption 2 we looked at the normalized change: $\langle\frac{\delta r_{e}(q)}{\lvert\lvert\delta r_{e}(q)\rvert\rvert},U^{T}e_{i}\rangle$. Here we will look at $\langle\delta r_{e}(q),U^{T}e_{i}\rangle$, so we expect the distribution to be:
$\langle\delta r_{e}(q),U^{T}e_{i}\rangle\sim\lvert\lvert\delta r_{e}(q)\rvert\rvert\cdot D$
Meaning the standard deviation scales linearly with the norm of $\delta r_{e}(q)$. Since $r_{e}$ scales linearly with $\delta r_{e}(q)$, we expect the standard deviation to also scale linearly with $r_{e}$. To measure the effective randomness, we look at $\langle\delta r_{e}(q),U^{T}(e_{i}-e_{correct})\rangle$, which shows explicitly that the correct answer logit change is sometimes enhanced and sometimes decreased relatively to the incorrect answers. We will observe that the noise is approximately normal.
To create the plot, for each question in a dataset, we look at the top 10 answers $e_{i},i\in[10]$ (with no representation engineering). We note that experimentally, the top 10 tokens make the majority of the probability mass (over 90%). Now for a given $r_{e}$ coefficient, we calculate the projection of these answers on $\delta r_{e}(q)$. We then aggregate these projections for all the questions in a few dataset and look at their histogram and at their standard deviation. We repeat this for different $r_{e}$ norms.
@@
The tangent of the curve of figure 10c,d is $\lambda\sigma$, as the curve is the standard deviation of $\langle\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|},U^{T}e_{i}\rangle\cdot|\delta r_{e}(q)|=\langle\frac{\delta r_{e}(q)}{|\delta r_{e}(q)|},U^{T}e_{i}\rangle\cdot\lambda r_{e}$, from assumption 3, and the inner product is a random variable of standard deviation $\sigma$, hence the tangent is $\lambda\sigma$. We observe that the noise is approximately normal. From the linear curve, we estimate $\lambda\sigma=0.5$, thus $\lambda\sigma\beta\approx 0.8\cdot 0.5$, as it is the mean of a half-normal distribution with parameter $\lambda\sigma$, which is approximately $0.8\lambda\sigma$.
@@
In figure 12 and 13, we plot the distance between the centers of representation clusters for positive and negative answers to harmful queries as the norm of harmful vectors is increased, for Llama 2 13B chat and Llama 2 13B respectively. As can be seen, the distance between the clusters increases, which corresponds to an increase in $\mathbb{E}[\langle\delta r_{e}(q),U^{T}(e_{good}-e_{bad})\rangle]$. We can define a range of coefficients in which the increase is bounded from below by a linear curve of the form in equation 11, meaning that the change in the model’s representation separates the positive and negative answer representations, similarly to the definition of $\Delta$-representation separability, but with mean instead of min. Thus by equation 11, the tangent of the lower bounding lines of figures 12 and 13 are an estimate for $\Delta\lambda$. From, this we get that $\Delta\lambda$ is approximately $0.1-0.3$